AC 223 Cost Accounting Hien Au & 10/17/18 922 PM Homework: Chapter 9 Homework Save core: 2.82 of 6 pts 3 of 5 14 complete) Hw Score: 48.68%,14.61 of 30 pe P9-37 (similar to) Question Hep ts product to Lucky Lager at $42 per banel Peter Bryant, Lucky Lagers coneroler, obtains the fellowing information Lucky Lager has just purchased the Pitsburgh Brewery. The brewery is two years old and uses absorption costng. ws about Pittsburgh Brewerys capacity and budgeted fxed manufacturing costs for 2017 EB (Click the lcon to view the information) Read the requirements
media%2Fffb%2Fffb78ce4-6353-4df3-a910-95
media%2Fd35%2Fd3556e1f-752f-4cc0-af33-c7
Answer #1

1.

Denominator-Level Capacity Concept Budgeted Fixed Manufacturing Overhead per Period Number of Barrels Budgeted Fixed Manufacturing Overhead Rate per Barrel
Theoretical Capacity $ 28,300,000 5,212,800 $ 5.43
Practical Capacity $ 28,300,000 3,500,000 8.09
Normal Capacity Utilization $ 28,300,000 2,835,000 9.98
Master Budget Capacity Utilization

$ 14,150,000

$ 14,150,000

2,835,000

9.98

The budgeted fixed manufacturing overhead rate is different in each case because the denominator is different for each type of capacity, while the numerator remains unchanged.

2.

Theoretical Capacity Practical Capacity Normal Capacity Utilization
Unit Sales 2,440,000 2,440,000 2,440,000
Sales Revenue $ 102,480,000 $ 102,480,000 $ 102,480,000
Less: Cost of Goods Sold
Variable Cost of Goods Sold 73,322,000 73,322,000 73,322,000
Fixed Manufacturing Overhead 13,249,200 19,739,600 24,351,200
Cost of Goods Sold ( Unadjusted 86,571,200 93,061,600 97,673,200
Adjustment for under/ over recovery of fixed manufacturing overhead 12,356,884 5,866,484 1,254,884
Cost of Goods Sold ( Adjusted) 98,928,084 98,928,084 98,928,084
Operating Income $ 3,551,916 $ 3,551,916 3,551,916

Unit sales = Beginning inventory + Production - Ending Inventory = 0 + 2,630,000 - 190,000 = 2,440,000 barrels.

Sales revenue = 2,440,000 x $ 42 = $ 102,480,000.

Variable cost of goods sold = ($ 79,031,500 / 2,630,000 ) * 2,440,000 = $ 73,322,000.